# Context Matters: Learning Generalizable Rewards via Calibrated Features

### v2

@@ 5.1 User Study
We conducted an in-person, IRB-approved study with 12 participants (7F/5M, ages 22-65, avg=29), each paid \$25 for 70-90 minute sessions. Participants taught their contextual preferences for two Utensil environment features, human distance and stove distance, (with discretized contexts for visualization purposes), with feature order counterbalanced. After an orientation on contextual feature queries and a familiarization round (50 queries for laptop distance from Cup environment), participants completed two experimental rounds (100 queries each). See App. D for study details.
@@
Manipulated variable. For each feature, we trained calibrated features using three different datasets: the first 25 queries, first 50, and all 100. Participants inspected and evaluated four unlabeled models using a graphical user interface: these calibrated feature variants and the uncalibrated base feature.
@@
Evaluation metrics. Participants completed a seven-point Likert scale survey, evaluating how well each model captured their personal contextual preference. The 4-item scale combined items adapted from Hoffman’s goal subscale hoffman2019evaluating (S1: The robot perceives accurately what my preferences are., S2: The robot does not understand how I would like the task to be completed.), prior work in the assessment of adaptive agents nikolaidis2017human (S3: The robot’s learned relationship is responsive to my input.), and a task-specific question (S4: The robot’s learned relationship corresponds well with my expectations for how it should behave in different contexts.). Participants also ranked the presented models based on how well each captured their contextual preferences.
@@
Hypothesis. Increasing the number of human-labeled contextual feature queries used in training improves alignment between the learned calibrated features and human preferences.
5.2 Results
We found partial support for our hypothesis: both Likert ratings and average rankings improved with the number of contextual feature queries used to train the calibrated feature (Fig. 5). A two-way repeated measures ANOVA (feature type × number of training queries) revealed a significant main effect of number of queries, $F(3,33)=19.49,p<0.001,\eta^{2}_{p}=0.427$, but no effect of feature type, $F(1,11)=0.09,p=0.77,\eta^{2}_{p}=0.0005$, and no interaction, $F(3,33)=1.33,p=0.28,\eta^{2}_{p}=0.0333$. Bonferroni-corrected post-hoc tests showed significant differences in Likert ratings between all calibrated features and the base feature $(p<0.05)$, and between 25 and 50 queries $(p<0.05)$. Both average Likert scores and average rankings (1 = best, 4 = worst) improved with the number of training queries, though values leveled off between 50 and 100 queries; see App. D for details.
Overall, calibrated features trained with any amount of contextual feature queries were significantly more aligned with user preferences than the base feature ($p<0.05$). Additionally, features trained with 50% of queries showed significantly better alignment than those trained with 25% ($p<0.05$). While further significance was not observed, likely due to high variance across only 12 participants, the subjective ratings trended positively with training query number. Thus, we find general support that increasing contextual feature queries improves alignment with user contextual preferences.
Finally, post-hoc analysis revealed diverse user preferences, which we grouped into shared high-level categories based on user descriptions. While we observe general trends of contextual preference, there is notable individual variation (Fig. 6). For the human distance feature, 6/12 participants preferred a similar general preference while the remaining 6 described 5 different variations of this preference. For the stove distance feature, 7/12 of participants preferred a similar general preference, while 5/12 held unique variants. These results suggest that while crowdsourcing or warm-starting general preferences may be useful, personalization is crucial to account for individual variations.
@@ 5.2 Results
Figure 6: Examples of two unique user contextual preferences for how utensil sharpness affects EE-human distance. Users successfully train calibrated features that accurately capture their preferences.
6 Conclusion
Discussion and broader impact. Our work demonstrates that explicit contextual modeling contributes to more flexible and efficient deployment of personalized robots in human environments. By targeting data to teach contextual effects, our method uses human input more effectively, producing representations that better generalize to new preferences without requiring multiple training preferences as demanded by multi-task baselines. The compositionality of our method allows calibrated features to be seamlessly incorporated in new tasks with new features without retraining—a modularity absent in multi-task baselines that lowers the barrier to efficient re-tasking of robots.
Our user study shows that real humans can effectively teach and validate calibrated features using contextual feature queries. This transparency supports iterative refinement and appropriate trust calibration, a critical component for responsible deployment of robots. In contrast, contextual effects remain hidden in the opaque representations of baseline methods. Finally, the diverse contextual preferences revealed in our user study further emphasize the need for sample-efficient learning to achieve preference alignment in widespread deployment.
Limitations. Our method faces several limitations. In continuous spaces, pairwise comparison learning with random sampling can fail to cover important regions of the input space. In Fig. 2 we see this issue result in a failure to capture feature irrelevance in specific contexts, such as when the stove is off or the cup is empty. This problem also affects scenarios where function variation only exists in small portions of the state space; when the cup is full, the calibrated laptop distance collapses to assign low values across most states, whereas the correct representation would assign low values except at the point cloud edge. We don’t observe the same failures in the discrete contexts of the user study (Fig. 6). To address this challenge, methods for improving sample efficiency and data spread could be explored, such as active learning techniques or data augmentation strategies. Future work could explore methods for gathering dense data within single important contexts, potentially drawing inspiration from feature traces bobu2022inducing . Finally, our method requires interpretable base features. Recent work has found ways for LLMs to generate reward features as code automatically from raw state peng2024adaptive , which is a complementary approach that could be used to generate a set of base reward features.
@@ Appendix A Implementation details
A.1 Loss regularization
We implemented the paired comparisons loss described in the main paper with an additional regularization term that penalizes the size of the model logits. For training the reward function $R_{\theta}$, this term regularization is,
$\ell_{\text{reg}}(\theta;\mathcal{D})=\sum_{(s_{1},s_{2},\ell)\in\mathcal{D}}(R_{\theta}(s_{1})^{2}+R_{\theta}(s_{2})^{2}).$ (5)
Similarly, for training calibrated feature function $\phi^{\prime}_{\psi_{i}}$, the analogous regularization term is,
$\ell_{\text{reg}}(\psi_{i};\mathcal{D})=\sum_{(s_{1},s_{2},\ell)\in\mathcal{D}}(\phi^{\prime}_{\psi_{i}}(s_{1})^{2}+\phi^{\prime}_{\psi_{i}}(s_{2})^{2}).$ (6)
@@ A.1 Loss regularization
$\mathcal{L}(\theta)=\frac{1}{|D_{\theta}|}\left(\lambda_{\text{RE}}\cdot\ell(\theta;D_{\theta}^{\text{equiv}})+\ell(\theta;D_{\theta}^{\text{pref}})+\lambda_{\text{reg}}\cdot\ell_{\text{reg}}(\theta;D_{\theta})\right).$ (7)
@@
$\mathcal{L}(\psi_{i})=\frac{1}{|D_{\psi_{i}}|}\left(\lambda_{\text{FE}}\cdot\ell(\psi_{i};D_{\psi_{i}}^{\text{equiv}})+\ell(\psi_{i};D_{\psi_{i}}^{\text{pref}})+\lambda_{\text{reg}}\cdot\ell_{\text{reg}}(\psi_{i};D_{\psi_{i}})\right).$ (8)
@@
A.2 Model architectures and hyperparameters
A.2.1 Selection
@@
A.2.2 Representation learning (pre-training)
@@
Table 1: Representation learning hyperparameters.
Calibrated Features Adam $10^{-3}$ 32 0.01 $10^{-4}$ 10 500
Joint/SinglePref Adam $10^{-3}$ 32 0.01 $10^{-4}$ 10 3000
A.2.3 Reward learning
Reward learning hyperparameters are listed in Table 2. For our method, we learned a linear reward over calibrated features, with frozen representations performing best. Additionally, our method performed the best skipping equivalence queries and generating new non-equivalent queries instead, referenced by the null $\lambda_{\text{FE}}$ value. All multi-task variants used a reward network consisting of a single 32-unit MLP layer (with PyTorch default weight initialization and standard ReLU activations). In low-data settings (5-10 preference queries), multi-task baselines with frozen and unfrozen representations performed similarly, but ultimately unfrozen representations performed best as more data was provided (see Fig. 15&16). Multi-task baseline models with unfrozen representations are the models visualized in the main paper figures.
Table 2: Reward learning hyperparameters.
Calibrated Features Adam $10^{-2}$ 32 0 0 - 200
Joint/SinglePref Adam $10^{-4}$ 64 $10^{-3}$ $10^{-3}$ 1 500
A.3 Compute resources
@@ Appendix B Environment details
We define three distinct environments, each characterized by three features as shown in blue in Fig. 7. The state for each environment is represented as a 23-dimensional continuous vector consisting of end-effector (EE) position (3 values: xyz coordinates) and orientation (9 values: rotation matrix), human position (3 values: xyz coordinates), stove position (3 values: xyz coordinates), laptop position (3 values: xyz coordinates), and 2 environment-specific contextual elements. We generate 10k random states per environment by selecting a random robot EE xyz location above the table within the robot’s reachable workspace, then solving inverse kinematics to obtain joint angles without collisions. The positions of other objects (human, stove, and laptop) remain constant across states. The two contextual elements of a state are randomly sampled with continuous values that range between 0 and 1, however our method can work with contextual elements that follow the state initialization generally in $\mathbb{R}$.
For each environment, all features are normalized to fall within the $[0,1]$ range using min-max normalization across the entire 10k-state dataset. For training stability, the state representation is similarly normalized.
In the Weighted Block environment, the features are z distance to table, xy-planar distance to stove, and xy-planar distance to laptop. These are affected by contextual elements of block weight and stove heat. This environment models scenarios where heavier objects may preferably be carried closer to the table and away from the laptop, while generally maintaining distance from a hot stove.
For the Cup environment, the features are cup angle, xy-planar distance to stove, and xy-planar distance to laptop. These are affected by contextual elements of cup fullness and stove heat. This environment captures situations where fuller cups should be carried more upright and away from the laptop, while generally maintaining distance from a hot stove.
In the Utensil environment, the features are utensil point at person, xy-planar distance to stove, and xy-planar distance to human. These are affected by contextual elements of utensil sharpness and stove heat. This environment represents scenarios where sharper utensils should be carried pointed away from and at a safe distance from the human, while generally maintaining distance from a hot stove.
Figure 7: Three experimental environments with features shown in blue.
The six unique feature definitions are detailed below:
  • Table distance measures the vertical distance between the EE and the table surface:
    $f_{\text{table}}(s)=p^{z}_{\text{EE}}-p^{z}_{\text{table}},$ (9)
@@
  • Laptop distance measures the planar (XY) distance between the EE and the laptop, capped at a maximum value,
    $f_{\text{laptop}}(s)=\min(\|p^{\text{XY}}_{\text{EE}}-p^{\text{XY}}_{\text{laptop}}\|_{2},0.8),$ (10)
@@
  • Human distance measures the planar (XY) distance between the EE and the human, capped at a maximum value,
    $f_{\text{human}}(s)=\min(\|p^{\text{XY}}_{\text{EE}}-p^{\text{XY}}_{\text{human}}\|_{2},1.0)$ (11)
@@
  • Point at human measures how much a utensil held by the robot points toward the human,
    $f_{\text{point}}(s)=\vec{u}^{W}_{\text{Ex}}{}^{\top}\vec{u}^{W}_{\text{EH}}$ (12)
@@
    $\vec{u}^{,W}_{\text{Ex}}=R^{W}_{\text{EE}}\begin{bmatrix}1,\ 0,\ 0\end{bmatrix}^{\top},\quad\vec{u}^{,W}_{\text{EH}}=\frac{p^{W}_{\text{human}}-p^{W}_{\text{EE}}}{\left\|p^{W}_{\text{human}}-p^{W}_{\text{EE}}\right\|_{2}}.$ (13)
    Here, $\vec{u}^{,W}_{\text{Ex}}$ is the x-axis of the end-effector (i.e., the utensil’s "pointing" direction) expressed in the world frame, obtained via the rotation matrix $R^{W}_{\text{EE}}$ from the end-effector to the world frame. The term $\vec{u}^{W}_{\text{EH}}$ describes the unit vector from the end-effector to the human, also in the world frame. The function equals $1$ when the utensil points directly at the human, and $-1$ when it points directly away.
@@
    $f_{\text{coffee}}(s)=\left[R^{W}_{WE}\right]_{3,1}$ (14)
    where $\left[R^{W}_{WE}\right]_{3,1}$ denotes the element in the third row and first column of the rotation matrix from the EE frame to the world frame. This represents the z-component of the EE’s x-axis in world coordinates. Since the cup’s upright direction is aligned with the EE’s x-axis, when this x-axis aligns with the world z-axis (vertical), the cup angle value approaches 1, indicating maximum "uprightness".
  • Stove distance The stove feature measures the planar (XY) distance between the EE and the stove, capped at a maximum value:
    $f_{\text{stove}}(s)=\min(\|p^{\text{XY}}_{\text{EE}}-p^{\text{XY}}_{\text{stove}}\|_{2},0.8)$ (15)
@@
Discrete contexts for visualization. While all simulated human participant experiments use continuous contexts, visualization requires establishing a set of discrete values for each contextual element (such as a discrete set of utensils varying in sharpness). Each contextual element is discretized with different granularities, all evenly spaced between 0 and 1: stove heat (8 discrete values), block weight (16 discrete values), cup fullness (6 values), and utensil sharpness (4 values). These discrete values are used only for the user study and generating visualization states for qualitative feature visualizations, as detailed in Appendix C.5.
Appendix C Simulated human experiments
C.1 Human simulation
C.1.1 Query response
Ground truth rewards used for simulating human responses are defined as linear combinations of the three environment ground truth calibrated features. In the scenario where only a single feature is contextually affected, two of the three calibrated features behave as identity functions, while the contextually affected feature is transformed according to the ground truth functions described in App. C.1.3. In the scenario where all three features are contextually affected, each is calibrated according to its respective ground truth calibrated feature function.
Ground truth linear reward weights, denoted $\theta^{*}$, are sampled from the set $\{-1,-0.5,0,0.5,1\}$, resulting in a total of 125 possible rewards per environment. A set of 10 test rewards is held out from training multi-task baselines and used solely for evaluation. The SinglePref baseline is trained on a single reward where all calibrated features are equally weighted: $\theta^{*}=[1,1,1]$. Each JointPref ($N$) model includes all rewards used in smaller variants; e.g., JointPref (50) contains the 25 rewards from JointPref (25) plus 25 additional ones. All variants include $\theta^{*}=[1,1,1]$ and $N-1$ additional training rewards.
Human preference selection is simulated using the Bradley-Terry model for pairwise comparisons [6]. Given two states, $s_{1}$ and $s_{2}$ the probability of selecting state $s_{1}$ is
$P_{\theta^{*}}(s_{1}\succ s_{2})=\frac{\exp(\beta\cdot R_{\theta^{*}}(s_{1}))}{\exp(\beta\cdot R_{\theta^{*}}(s_{1}))+\exp(\beta\cdot R_{\theta^{*}}(s_{2}))}\quad,$ (16)
where $\beta=20$ is chosen based on empirical validation. The same setup is used to simulate responses to contextual feature queries, except $\phi^{\prime}_{i}$ is used in place of $R_{\theta^{*}}$ for the $i$-th calibrated feature. The value of $\beta$ is kept constant across all query types. Note that this is the same model of human preference selection used during learning (Eq. 1), although during learning we assume $\beta$ is unknown and fix $\beta=1$.
To simulate equivalence selections, we define a threshold such that two states are marked as equivalent if the absolute difference in their ground truth value (either reward or a specific calibrated feature) falls within that threshold. This threshold was empirically determined, as described in App. C.1.2.
Figure 8: Equivalence threshold experiments for calibrated features.
Figure 9: Equivalence threshold experiments for multi-task baselines.
C.1.2 Equivalence threshold selection
To simulate human judgments of equivalence in paired comparisons, we introduce a threshold $\epsilon$ that defines when two states are considered equivalent. Specifically, this threshold determines whether $|f(s_{1})-f(s_{2})|\leq\epsilon$, in which case the simulated human labels the pair as equivalent, i.e., $f(s_{1})=f(s_{2})$. The function $f$ represents any operation on the state, such as a calibrated feature value or the reward. Using a common $\epsilon$ across both types is reasonable, since the ground truth reward is simply a linear combination (with weights with maximum magnitude of 1), of three calibrated features (each with maximum magnitude of 1).

### v3

@@ 5. User Study
We conducted an in-person, IRB-approved study with 12 participants (7F/5M, aged 22-65, avg=29) recruited from the campus community, each paid \$25 for 70-90 minute sessions. Participants taught their contextual preferences for two Utensil environment features, human and stove, with feature order counterbalanced and contexts discretized for visualization purposes. After instruction on contextual feature queries and a familiarization round (50 queries for laptop feature in Cup environment), participants completed the two experimental rounds (100 queries each). More details about the surveys and user interface can be found in App. D.1D.2.
@@
Manipulated variable. For each feature, we trained calibrated features using three different datasets: the first 25 queries, first 50, and all 100. Participants used a graphical user interface to inspect and evaluate four unlabeled models: these calibrated feature variants and the uncalibrated base feature.
@@
Dependent variables. Participants first wrote a brief description of their contextual preference, then responded to 100 queries with response times recorded. After query response, participants completed the System Usability Scale (SUS) and NASA-TLX. They then inspected and subjectively assessed the four presented models on a seven-point Likert scale, evaluating how well each captured their contextual preferences. The 4-item scale combined items adapted from Hoffman’s goal subscale (Hoffman, 2019) (S1: The robot perceives accurately what my preferences are., S2: The robot does not understand how I would like the task to be completed.), prior work in the assessment of adaptive agents (Nikolaidis et al., 2017) (S3: The robot’s learned relationship is responsive to my input.), and a task-specific question (S4: The robot’s learned relationship corresponds well with my expectations for how it should behave in different contexts.). Participants then ranked the models by how well each captured their contextual preferences.
@@
Hypothesis. Training with more user-labeled contextual feature queries improves alignment between learned calibrated features and the user’s contextual preferences.
Examples of two different user contextual preferences for the utensil feature, depicted by the user’s written preference description the learned calibrated feature. Learned calibrated features are evaluated at all states in a point cloud and shown at different levels of utensil sharpness.
@@
Figure 6. Examples of two unique user contextual preferences for how utensil sharpness should affect the human feature ($E$ xy-planar distance to human). Users differ in their comfort with robot proximity and in which utensils they view as similar. By labeling contextual feature queries, users trained calibrated features that accurately captured these preferences.
Examples of two different user contextual preferences for the utensil feature, depicted by the user’s written preference description the learned calibrated feature. Learned calibrated features are evaluated at all states in a point cloud and shown at different levels of utensil sharpness.
5.1. Results
The results predominantly support our hypothesis: both Likert ratings and average rankings improved with the number of contextual feature queries (Fig. 5). A two-way repeated measures ANOVA (feature type × query count) revealed a significant main effect of query count, $F(3,33)=19.49,p<0.001,\eta^{2}_{p}=0.427$, but no effect of feature type, $F(1,11)=0.09,p=0.77,\eta^{2}_{p}=0.0005$, and no interaction, $F(3,33)=1.33,p=0.28,\eta^{2}_{p}=0.0333$. Bonferroni-corrected post-hoc tests showed significant differences in Likert ratings between all calibrated features and the base feature $(p<0.05)$, and between 25 and 50 queries $(p<0.05)$. Average Likert scores and rankings (1 = best, 4 = worst) improved with the number of training queries, though values leveled off between 50 and 100 queries.
Overall, calibrated features trained with any amount of contextual feature queries were significantly more aligned with user preferences than the base feature ($p<0.05$). Additionally, features trained with 50% of queries showed significantly better alignment than those trained with 25% ($p<0.05$). While further significance was not observed, likely due to variance across 12 participants, subjective ratings consistently increased with more queries. Thus, we find general support that increasing contextual feature queries improves alignment with user contextual preferences.
The user study additionally demonstrated the practical feasibility of the proposed method. The average response time per feature (100 queries) was 11.56 $\pm$ 2.36 minutes, indicating that contextual feature queries provide an efficient way to capture these personal contextual preferences. Moreover, Fig. 5 shows different features need different amounts of data to converge in training, ranging from 50 to 100 queries in our experiments. The stove feature showed similar Likert scores at 50 and 100 queries, indicating 50 queries sufficed. Participants also reported acceptable usability on the SUS (M = 71.0, SD = 18.3) and moderate workload on the NASA-TLX: mental (8.0 ± 5.1), physical (1.8 ± 1.9), temporal (9.6 ± 4.2), performance (13.8 ± 3.3), effort (8.0 ± 4.9), and frustration (4.1 ± 3.6), with both metrics averaged over the two query response rounds.
Finally, post-hoc analysis of written user preference descriptions revealed diverse preferences, which we grouped into shared high-level categories (see App. D.3 for analysis details and all descriptions). While we observe general trends of contextual preference, there is notable individual variation (Fig. 6). For the human feature, 6/12 participants preferred a similar general preference while the remaining 6 described 5 different variations. For the stove feature, 7/12 of participants preferred a similar general preference, while 5/12 held unique variants. These results suggest that while crowdsourcing or warm-starting general preferences may be useful, personalization is crucial to account for individual variations.
6. Discussion and Limitations
Our method faces several limitations. In continuous spaces, randomly sampled pairwise comparisons can miss important regions of the state space, which hurts our method’s sample efficiency. Fig. 2 shows this omission can prevent capturing feature irrelevance in specific contexts, such as when the stove is off or cup empty. It can also miss variations in small regions of the state. For example, when the cup is full, the calibrated laptop feature collapses to low values across most states, although the point-cloud edge should remain high. These issues do not appear in the user study discrete contexts (Fig. 6). Future work could improve sample efficiency and coverage through active learning or dense data collection in key contexts, drawing on feature traces (Bobu et al., 2022). Also, our post hoc analysis (App. D.3) suggests users’ contextual preferences often share similarities, implying representations could be warm-started with crowdsourced data and refined via our method for necessary personalization.
Additionally, our method requires interpretable base features. Recent works use large-language or vision-language models to generate reward features as code automatically from raw state (Chen et al., 2025; Peng et al., 2024), which are complementary approaches that could be used to generate a set of base reward features. In our user study we did not test whether users can identify which base features are contextually-affected, but their semantic meaning suggests this could be feasible in practice. Prior work shows that users naturally reference and correct individual features when behavior is inappropriate (Bajcsy et al., 2017). Thus, we hypothesize that in deployment contextualized features can be obtained by either (1) asking users which features should vary with context, or (2) inferring the implicated feature from user corrections.
Finally, we consider practical limits to scaling to larger state spaces (e.g., images). Future work could explore composing state representations from image features (e.g., using DINOv3 (Siméoni et al., 2025)) and improving sample efficiency through active learning. When environments change (e.g., new objects are added), existing calibrated features remain valid unless the input state space changes. However, while new base features (e.g., distances to new objects) may share similar contextual adaptations (e.g., further EE distance depending on manipulated cup fullness), our current pipeline requires a new calibrated feature to be learned for each base feature. Future work could explore learning representations that generalize across object types, such as all “technology” or all “fragile” objects.
In summary, we show that explicit contextual modeling enables more efficient learning of robot rewards that adapt across contexts. By targeting data to teach contextual effects, our method learns useful representations from minimal human input. Its feature-based structure allows learned representations to be modularly composed into new rewards–a capability missing in baselines. This structure also suggests representations have potential for interpretability. Future work could assess interpretability for supporting iterative refinement and appropriate trust calibration, both essential for responsible robot deployment. Our user study shows people can teach their contextual preferences, and the diversity of these preferences underscores the value of our method’s sample efficiency. By making it easier for users to teach complex behaviors, we take a step toward achieving preference alignment in widespread robot deployment.
Acknowledgements.
We thank Mike Hagenow and Ho Chit Siu for helpful discussions. This work was partially funded by the Department of Defense (DoD) through the National Defense Science and Engineering Graduate (NDSEG) Fellowship Program.
@@ Appendix A Implementation details
Table 1. Representation learning hyperparameters.
Calibrated Features Adam $10^{-3}$ 32 0.01 $10^{-4}$ 10 500
Joint/SinglePref Adam $10^{-3}$ 32 0.01 $10^{-4}$ 10 3000
Table 2. Reward learning hyperparameters.
Calibrated Features Adam $10^{-2}$ 32 0 0 - 200
Joint/SinglePref Adam $10^{-4}$ 64 $10^{-3}$ $10^{-3}$ 1 500
RandomInit Adam $10^{-3}$ 32 $10^{-2}$ 0 1 500
A.1. Loss regularization
We implemented the paired comparisons loss (Eq. 45) with an additional regularization term that penalizes the size of the model logits. For training the reward function $R_{\theta}$, this term regularization is,
(6) $\ell_{\text{reg}}(\theta;\mathcal{D})=\sum_{(s_{1},s_{2},\ell)\in\mathcal{D}}(R_{\theta}(s_{1})^{2}+R_{\theta}(s_{2})^{2}).$
Similarly, for training calibrated feature function $\phi^{c}_{\psi_{i}}$, the analogous regularization term is,
(7) $\ell_{\text{reg}}(\psi_{i};\mathcal{D})=\sum_{(s_{1},s_{2},\ell)\in\mathcal{D}}(\phi^{c}_{\psi_{i}}(s_{1})^{2}+\phi^{c}_{\psi_{i}}(s_{2})^{2}).$
@@ A.1. Loss regularization
(8) $\mathcal{L}(\theta)=\frac{1}{\lvert D_{\theta}\rvert}\left(\lambda_{\text{RE}}\cdot\ell(\theta;D_{\theta}^{\text{equiv}})+\ell(\theta;D_{\theta}^{\text{pref}})+\lambda_{\text{reg}}\cdot\ell_{\text{reg}}(\theta;D_{\theta})\right).$
@@
(9) $\mathcal{L}(\psi_{i})=\frac{1}{|D_{\psi_{i}}\rvert}\left(\lambda_{\text{FE}}\cdot\ell(\psi_{i};D_{\psi_{i}}^{\text{equiv}})+\ell(\psi_{i};D_{\psi_{i}}^{\text{pref}})+\lambda_{\text{reg}}\cdot\ell_{\text{reg}}(\psi_{i};D_{\psi_{i}})\right).$
@@
A.2. Model architectures and hyperparameters
Here we present the selected architecture and hyperparameters and describe the selection process. For training stability, model inputs (state and features) were normalized to the $[0,1]$ range across the full 10k-state dataset in an environment.
A.2.1. Selection
@@
A.2.2. Representation learning (pre-training)
@@
A.2.3. Reward learning
Reward learning hyperparameters are listed in Table 2. For all methods we evaluated freezing or fine-tuning pre-trained representations, with details in App. C.3. For our method, we learned a linear reward over calibrated features. Our method performed the best skipping equivalence queries and generating new non-equivalent queries instead, referenced by the null $\lambda_{\text{FE}}$ value. All multi-task variants used a reward network consisting of a single 32-unit MLP layer (with PyTorch default weight initialization and standard ReLU activations). For the RandomInit baseline, we tested learning rewards from scratch using the same end-to-end architecture as JointPref, and with no compressed embedding, finding the same architecture as JointPref to perform the best. We performed a separate round of hyperparameter tuning for RandomInit, with selected hyperparameters also in Table 2.
A.3. Compute resources
@@
Three manipulation environments. The robot is on a table near a stove and a laptop, and a human is standing nearby. The robot holds either a block, a cup, or a utensil. Features are visually depicted using blue arrows and labeled with text.
Figure 7. Three experimental environments with features shown in blue.
Three manipulation environments. The robot is on a table near a stove and a laptop, and a human is standing nearby. The robot holds either a block, a cup, or a utensil. Features are visually depicted using blue arrows and labeled with text.
@@ Appendix B Environment details
We generate 10k random states per environment by selecting a random robot EE xyz location above the table within the robot’s reachable workspace, then solving inverse kinematics to obtain joint angles without collisions. Object positions (human, stove, and laptop) remain constant across states. The two contextual elements of a state are randomly sampled with continuous values that range between 0 and 1; however, our method can work with contextual elements that follow the state initialization generally in $\mathbb{R}$.
B.1. Base feature definition
Fig. 7 depicts the three base features in each manipulation environment. The six unique base features definitions are detailed below:
  • Table measures the vertical distance between the EE and the table surface:
    (10) $f_{\text{table}}(s)=p^{z}_{\text{EE}}-p^{z}_{\text{table}},$
@@ B.1. Base feature definition
  • Laptop measures the planar (XY) distance between the EE and the laptop, capped at a maximum value,
    (11) $f_{\text{laptop}}(s)=\min(\|p^{\text{XY}}_{\text{EE}}-p^{\text{XY}}_{\text{laptop}}\|_{2},0.8),$
@@
  • Human measures the planar (XY) distance between the EE and the human, capped at a maximum value,
    (12) $f_{\text{human}}(s)=\min(\|p^{\text{XY}}_{\text{EE}}-p^{\text{XY}}_{\text{human}}\|_{2},1.0)$
@@
  • Point measures how much a utensil held by the robot points toward the human,
    (13) $f_{\text{point}}(s)=\vec{u}^{W}_{\text{Ex}}{}^{\top}\vec{u}^{W}_{\text{EH}}$
@@
    (14) $\vec{u}^{,W}_{\text{Ex}}=R^{W}_{\text{EE}}\begin{bmatrix}1,\ 0,\ 0\end{bmatrix}^{\top},\quad\vec{u}^{,W}_{\text{EH}}=\frac{p^{W}_{\text{human}}-p^{W}_{\text{EE}}}{\left\|p^{W}_{\text{human}}-p^{W}_{\text{EE}}\right\|_{2}}.$
    Here, $\vec{u}^{,W}_{\text{Ex}}$ is the x-axis of the end-effector (i.e., the utensil’s ”pointing” direction) expressed in the world frame, obtained via the rotation matrix $R^{W}_{\text{EE}}$ from the end-effector to the world frame. The term $\vec{u}^{W}_{\text{EH}}$ describes the unit vector from the end-effector to the human, also in the world frame. The function equals $1$ when the utensil points directly at the human, and 1-1 when it points directly away.
@@
    (15) $f_{\text{coffee}}(s)=\left[R^{W}_{WE}\right]_{3,1}$
    where $\left[R^{W}_{WE}\right]_{3,1}$ denotes the element in the third row and first column of the rotation matrix from the EE frame to the world frame. This represents the z-component of the EE’s x-axis in world coordinates. Since the cup’s upright direction is aligned with the EE’s x-axis, when this x-axis aligns with the world z-axis (vertical), the cup angle value approaches 1, indicating maximum ”uprightness”.
  • Stove The stove feature measures the planar (XY) distance between the EE and the stove, capped at a maximum value:
    (16) $f_{\text{stove}}(s)=\min(\|p^{\text{XY}}_{\text{EE}}-p^{\text{XY}}_{\text{stove}}\|_{2},0.8)$
@@
B.2. Discrete contexts for visualization.
While all simulated human participant experiments use continuous contexts, visualization requires establishing a set of discrete values for each contextual element (such as a discrete set of utensils varying in sharpness). Each contextual element is discretized with different granularities, all evenly spaced between 0 and 1: stove heat (8 discrete values), block weight (16 discrete values), cup fullness (6 values), and utensil sharpness (4 values). These discrete values are used only for the user study and generating visualization states for qualitative feature visualizations.
Appendix C Simulated Experiments
C.1. Human simulation
C.1.1. Equivalence threshold selection
To simulate human judgments of equivalence in paired comparisons, we introduce a threshold $\epsilon$ that defines when two states are considered equivalent. Specifically, this threshold determines whether $\lvert f(s_{1})-f(s_{2})\rvert\leq\epsilon$, in which case the simulated human labels the pair as equivalent, i.e., $f(s_{1})=f(s_{2})$. The function $f$ represents any operation on the state, such as a calibrated feature value or the reward. Using a common $\epsilon$ across both types is reasonable, since the ground truth reward is simply a linear combination (with weights with maximum magnitude of 1), of three calibrated features (each with maximum magnitude of 1).